OpenStudy (anonymous):

can anyone please help me prove this norm inequality ||x||_1 ≤ n*||x||_\infty, where ||x||_\infty = max |x_i|

OpenStudy (anonymous):

||x||_1 = x_1 + x_2 + ... + x_n ||x||_\infty = max(1<i<n) |x_i|

OpenStudy (anonymous):

ouch...its hard to make any of that out.